### v1

# Probing the axion-photon coupling with space-based gravitational wave detectors

Abstract
We propose a simple modification of space-based gravitational wave (GW) detector optical benches which would enable the measurement of vacuum birefringence of light induced by axion dark matter through its coupling to electromagnetism. While marginally affecting the sensitivity to GW by a factor $\sqrt{2}$, we show that such an adjustment would make future detectors such as LISA, TianQin, Taiji and Big-Bang Observer the most sensitive experiments at low axion masses.
I Introduction
Almost $100$ years after the first hints for dark matter (DM) through its gravitational effect on visible matter Zwicky (1933), its microscopic nature remains unknown and constitutes one of the biggest mysteries in modern physics Bertone and Tait (2018). Several cosmological and astrophysical observations (see e.g. Battaglieri et al. (2017); Carr et al. (2016)) inform us on the bounds of the mass of the DM constituent, ranging from $10^{-22}$ eV to $10^{68}$ eV. These $\sim 90$ orders of magnitude of mass imply that the search for the nature of DM constitutes an experimental challenge. Out of all the DM candidates, ultralight dark matter (ULDM), with masses below $1$ eV has recently gained a lot of attention because of lack of evidence of historically dominant models such as weakly interacting massive particles. In addition, the recent extraordinary development of low-energy precision measurement techniques Safronova et al. (2018) provides a hope of near-future detection for such low-mass candidates. One of the best-motivated models of ULDM is the axionlike-particle (in the following referred to as the axion), a pseudo-scalar DM candidate Weinberg (1978); Wilczek (1978), which in some circumstances, can also solve the strong CP problem Peccei and Quinn (1977). The coupling between such axion and electromagnetism (EM) arises naturally in many theories (see e.g. Sikivie (2021); Arias et al. (2012); Svrcek and Witten (2006)) and produces many different observable effects in many different systems (see Sikivie (2021) for a review). One of them is the phase velocity difference between the two circularly polarized states of light, known as vacuum birefringence Obata et al. (2018). In such a case, one could observe a phase shift between two circularly polarized beams with different polarizations or between a circularly polarized beam and a linearly polarized beam, whose dispersion relation is unchanged at first order.
Searches of the axion-photon coupling through vacuum birefringence have recently been performed using an optical cavity Michimura et al. (2020). In addition, a modification to ground-based gravitational wave (GW) detectors was proposed in order to be sensitive to such an effect Nagano et al. (2019). In this paper, we propose a new scheme to search for the axion, based on the interferometric optical system used for future space-based GW detectors, such as LISA Colpi et al. (2024), TianQin Luo et al. (2016), Taiji Luo et al. (2020) and Big-Bang Observer (BBO) Crowder and Cornish (2005). The principle of such space-missions relies on the measurement of small variations of the optical pathlength between free falling test-masses contained in distant spacecrafts (S/C). Currently, the standard design of optical benches (OB) of those experiments makes use only of linearly polarized light. With such a setup, these space-based GW detectors are insensitive to vacuum birefringence impacting circularly polarized light such as the one induced by axions. In this paper, we propose a slight modification to the design of such OB which would change the polarization of the laser beams exchanged between the spacecraft from linear to circular. With such a setup, the space-based GW observatory becomes sensitive to the axion-photon coupling, without jeopardizing the primary goal of the experiment, i.e detecting GW. More particularly, the proposed modification offers sensitivity to both GW and DM, whilst still satisfying the requirement that stray light from spurious reflections of the emission be rejected due to its polarization. Assuming this minor setup modification, we estimate the sensitivity of those various experiments to the axion-photon coupling, and we show that they would all reach unprecedented sensitivities, compared to existing bounds from astrophysics and cosmology, in a mass region where laboratory experiments are poorly represented. Together with Nagano et al. (2019), this study completes the potential search of the axion-photon coupling through future GW detectors, and we show that ground-based and space-based detectors would be complementary for this matter.
The paper is organised as follows. In Sec. II, we quickly derive the phase velocity difference between left and right circular polarized light induced by the axion-photon coupling. In Sec. III, we discuss the various requirements for GW and axion detection in terms of light polarization inside the S/C OB, and we show that a slight modification of the OB current design would make both detections possible. In Sec. IV, we compute the single-arm Doppler shift induced by the oscillation of the phase velocity of circularly polarized light, and we derive the sensitivity curves of each GW space-based detector, by making use of the appropriate Time Delay Interferometry (TDI) combination Armstrong et al. (1999); Tinto and Armstrong (1999); Dhurandhar et al. (2002). Finally, we provide a short discussion of our results.
II Vacuum birefringence
In this section, we derive the vacuum birefringence arising from the axion-photon coupling with strength $g_{a\gamma}$ (in units of GeV-1). We start from the action of the dimensionless axion pseudo-scalar field $a$ with mass $m_{a}$ that we define as
$S_{a}$ $=\frac{1}{c}\int d^{4}x\sqrt{-g}\left[\frac{R}{2\kappa}-\frac{1}{2\kappa}g^{\mu\nu}\partial_{\mu}a\partial_{\nu}a-\frac{m^{2}_{a}c^{2}}{2\hbar^{2}\kappa}\right]\,,$ (1)
where $R$ is the Ricci scalar, $g_{\mu\nu}$ is the spacetime metric tensor (with ($+$)(-\>+\>+\>+)( - + + + ) signature) with determinant $g$ and $\kappa=8\pi G/c^{4}$ is the Einstein gravitational constant. At cosmological scales, when $m_{a}c^{2}/\hbar\gg H$, with $H$ the Hubble constant, the axion field equation derived from Eq. (1) admits plane wave solutions, i.e.
$a(t,\vec{x})$ $=a_{0}\cos(\omega_{a}t-\vec{k}_{a}\cdot\vec{x}+\Phi)\,,$ (2a)
where $a_{0},\Phi$ are respectively the dimensionless amplitude and an unknown phase, and $|\vec{k}_{a}|=\sqrt{\omega^{2}_{a}/c^{2}-m^{2}_{a}c^{2}/\hbar^{2}}$ is the wavevector of the field. If the axion field is identified as the nonrelativistic DM, one has
$a_{0}$ $=\frac{\sqrt{16\pi G\rho_{\mathrm{DM}}}}{\omega_{a}c}\,$ (2b)
$\vec{k}_{a}$ $=-\frac{\omega_{a}v_{\mathrm{DM}}}{c^{2}}\hat{e}_{v}\,,$ (2c)
where $\rho_{\mathrm{DM}}=0.4$ GeV/cm3 is the local DM energy density McMillan (2011), $\vec{v}_{\mathrm{DM}}=v_{\mathrm{DM}}\hat{e}_{v}\sim 10^{-3}\>c\hat{e}_{v}$ is the mean velocity of the Sun in the galactic halo and
$\omega_{a}\sim\frac{m_{a}c^{2}}{\hbar}$ (2d)
(at leading order in $(v_{\mathrm{DM}}/c)^{2}$).
As we shall see in the following, we will be interested in experiments with typical length scales corresponding to the distance between the S/C exchanging light signals $L\leq 3\times 10^{9}$ m, sensitive to frequencies in the band $f_{a}\in[10^{-4},10]$ Hz. Thus, the de Broglie wavelength of the field exceeds largely the distance between the S/C $\lambda_{a}=2\pi/|\vec{k}_{a}|\geq 3\times 10^{10}\>\mathrm{m}\gg L$, or in other words, the axion field is highly homogeneous on the scale of the experiment. Therefore, the propagation phase of the plane wave can be neglected such that Eq. (2a) becomes
$a(t,\vec{x})$ $\approx\frac{\sqrt{16\pi G\rho_{\mathrm{DM}}}}{\omega_{a}c^{2}}\cos(\omega_{a}t+\Phi)\,,$ (3)
In addition to the action Eq. (1), we add the Lagrangian describing the interaction in vacuum between the axion $a$ and electromagnetism (EM) through the coupling $g_{a\gamma}$ Sikivie (2021)
$\mathcal{L}$ $=-\frac{1}{4\mu_{0}}F_{\mu\nu}F^{\mu\nu}-E_{P}\frac{g_{a\gamma}}{4\mu_{0}}aF_{\mu\nu}\tilde{F}^{\mu\nu}\,,$ (4)
where the EM gauge field $A^{\alpha}$ has usual units of V.s/m, where $F^{\mu\nu}$ is the EM strength tensor, $\tilde{F}^{\mu\nu}=1/2\epsilon^{\mu\nu\rho\sigma}F_{\rho\sigma}\equiv\epsilon^{\mu\nu\rho\sigma}\partial_{\rho}A_{\sigma}$ is the dual EM strength tensor, with $\epsilon^{\mu\nu\rho\sigma}$ the antisymmetric 4-dimensional Levi-Civita symbol, $\mu_{0}$ is the vacuum permeability and $E_{P}=\hbar c/\kappa\sim 10^{18}$ GeV is the reduced Planck energy.
From Eq. (4), we can derive the EM field equations
$\square{A^{\nu}}-\partial^{\nu}\partial_{\mu}A^{\mu}+E_{P}g_{a\gamma}(\partial_{\mu}a)\epsilon^{\mu\nu\rho\sigma}\partial_{\rho}A_{\sigma}$ $=0\,.$ (5)
We now focus on the spatial equation of motion of the EM field (i.e $\nu\neq 0$ in Eq. (5)). Assuming both temporal and Coulomb gauges for the EM field $A_{0}=0$ and $\vec{\nabla}\cdot\vec{A}=0$, one finds Obata et al. (2018)
$\ddot{A_{i}}-c^{2}\nabla^{2}A_{i}+cE_{P}g_{a\gamma}\dot{a}\epsilon_{ijk}\partial_{j}A_{k}=0\,.$ (6)
Decomposing light in the two right $A^{+}$ and left $A^{-}$ circular polarizations, Eq. (6) becomes Obata et al. (2018)
$\ddot{A}^{+}_{i}=-k^{2}c^{2}\left(1+\frac{\sqrt{16\pi G\rho_{\mathrm{DM}}}g_{a\gamma}E_{P}}{kc^{2}}\sin(\omega_{a}t+\Phi)\right)A^{+}_{i}\,$ (7a)
$\ddot{A}^{-}_{i}=-k^{2}c^{2}\left(1-\frac{\sqrt{16\pi G\rho_{\mathrm{DM}}}g_{a\gamma}E_{P}}{kc^{2}}\sin(\omega_{a}t+\Phi)\right)A^{-}_{i}\,,$ (7b)
respectively for the right and left polarizations. With a plane wave Ansatz, one finds that the left and right modes have different dispersion relations, i.e both circular polarizations travel with different phase velocities $c_{\pm}=\omega_{\pm}/k$, respectively Obata et al. (2018)
$c_{\pm}$ $=c\sqrt{1\pm\frac{\sqrt{16\pi G\rho_{\mathrm{DM}}}g_{a\gamma}E_{P}}{kc^{2}}\sin(\omega_{a}t+\Phi)}$ (8a)
$\equiv c\pm\delta c(t)\,,$ (8b)
with
$\delta c(t)\approx\frac{\sqrt{4\pi G\rho_{\mathrm{DM}}}g_{a\gamma}}{kc^{2}}\sin(\omega_{a}t+\Phi)\,.$ (8c)
In other words, vacuum becomes birefringent in presence of an axion background. Experiments such as DANCE Obata et al. (2018); Michimura et al. (2020) attempt to detect the axion-photon coupling through this effect. We will now show how space-based GW detectors could probe this coupling via a slight modification of their OB.
III Optical benches
In this section, we will be interested in the polarization of light inside the detectors OB and in between the S/C. As a reminder, all the experiments under consideration in this paper will consist of one (or more) constellation of three S/C, in the following noted $1,2,3$. Each S/C is hosting two OB (labeled $A$ and $B$) with one laser per OB. These two laser beams are sent to the two other S/C and GW signals will be detected by interferometric measurement between those optical signals (see Fig. 1). For our purposes, we are interested in the polarization of these beams inside the S/C and also in between the S/C.
Figure 1: The constellation involves three S/C, denoted $1,2,3$, each containing two optical benches as described in Fig. 2 denoted by ”A” and ”B”. Six different light signals are sent between the different S/C to form the interferometer; three vertically polarized beams are propagating clockwise and the other three horizontally polarized beams are propagating counterclockwise.
III.1 Current design for the optical benches
The current (simplified) design for the OB envisioned for space-based GW observatory is shown in Fig. 2, from Bayle and Hartwig (2023); Luo et al. (2016, 2020), where we only take into account the interference between laser beams of the local S/C with a distant one. We assume for simplicity that the light wave is produced with a linear polarization along the $x$-axis, is travelling along the $z$ direction and arrives on the detector with a polarization along the $y$-axis after passing through the half-wave plate.
III.2 Requirements for GW detection
Figure 2: Simplified OB design currently envisioned for space-based GW detectors. The two OB $A$ and $B$ are inside two different S/C (shown as the light green cube) and exchange laser beams between each other. The optical elements are shown in blue, the light polarization in red and the direction of propagation with black arrows. Inside each of these OB, an initial linear polarization directed along the $x$-axis is produced by the laser and the association of the optical elements inside each bench fulfill all the necessary requirements to enable GW detection (see text).
In terms of light polarization, several requirements are necessary for the detection of GW in space-based detectors, as follows
  • In each OB, the incoming beam (i.e the one which has travelled through free space from the distant S/C) must be, at least partially, reflected by the polarized beam splitter to enter the interferometer and interfere with the local beam.
  • When leaving one S/C, a small portion of the local beam will be undesirably reflected on it, come back inside the OB and create an additional noise in the interferometer [^1] Assuming the frequency of the local beam to be $f_{\ell}$, the distance beam has a slightly different frequency $f_{\ell}+\delta f$, due to standard Doppler from S/C differential velocity, and frequency planning, see Heinzel et al. (2024). The reflected component of the local beam could interfere with itself but would only create a DC component. However, it could also interfere with the distant beam, and create a spurious beatnote at $\delta f$. In addition, the motion of the reflecting component at the edge of the S/C would induce a time dependent phase on the reflected beam, which would mimic a signal from the GW, and greatly alter the measurement.. We require this reflected component to be rejected by the polarized beam splitter, such that it does not enter the interferometer.
It can be shown easily that the current OB design from Fig. 2 fulfills those requirements. In the full constellation (with the three S/C) with two OB in each S/C, six individual interference measurements are performed at each time. Two different linear polarizations travel in space : vertical polarization travels in one direction (say clockwise) while horizontal polarization travels in the other direction (say counterclockwise) (see Fig. 1).
III.3 Requirements for axion detection
In addition to the previous requirements for space-based GW detectors to work, the additional condition necessary for the axion-photon detection is
  • The two beams travelling in free space between the S/C (e.g. $A\rightarrow B$ and $B\rightarrow A$) must have a different polarization, and at least one of them has to be circularly polarized.
Indeed, as derived in Section. II, the axion field only impacts the phase velocity of circularly polarized light and not linearly polarized one [^2] In reality, the coupling also affects the two linear polarization states because it couples them, i.e a linear state slightly oscillates into the other one and vice-versa. Using e.g. an initial pure ”horizontal” polarization state, one could see such effect using a vertical polarizer and counting the photons that go through due to the rotation of the polarization plane. This is essentially the detection idea in e.g. Nagano et al. (2019).. Therefore, in order to detect the axion field one must introduce other optical elements inside the original OB shown in Fig. 2 in order to generate circularly polarized light, whilst still satisfying the requirements for GW detection of Sect. III.2.
Figure 3: Modified OB design for the OB $A$ (to be compared to the original one presented in Fig. 2), in order to be sensitive to vacuum birefringence induced by the axion-photon coupling. The half-wave plate at the end of the OB $A$ is replaced by a right polarizer which is a combination of a quarter-wave plate, a horizontal polarizer and a second quarter-wave plate, are oriented conveniently. This configuration produces a right polarized light at the output of OB $A$, and fulfill all the requirements for the optical interferometer to work. The OB $B$ is unchanged.
III.4 Proposed modification
A solution to all the requirements presented in the last two sections is to replace the last $\hat{\lambda}_{2}$ in the OB $A$ (see Fig. 2), by a $\hat{\lambda}_{4}(\pi/4)\hat{C}_{H}\hat{\lambda}_{4}(-\pi/4)$ combination, i.e a first quarter-wave plate oriented with $\theta=\pi/4$, then a horizontal linear polarizer and then a second quarter wave plate oriented with $\theta=-\pi/4$, as shown in Fig. 3. This corresponds to a right polarizer. In Appendix. A, we show mathematically how this solves the various requirements, using the Jones formalism. Therefore, through this modification of the OB, half of the beams are circularly polarized (see the modified full constellation in Fig. 4), and therefore the detection of the axion-photon coupling would be possible.
Figure 4: Proposed modified constellation for the search of vacuum birefringence, as a consequence of Fig. 3. In this framework, only the arms where light is circularly polarized are impacted by the axion field, and get phase shifted (see text).
IV Sensitivity of space-based GW observatory to the axion-photon coupling
In the previous section, we have demonstrated that a minor change in the OB design from space-baed GW observatories can make them sensitive to the coupling between an axion and electromagnetism. Let us now assess quantitatively the sensitivity of these space missions to $g_{a\gamma}$.
IV.1 Single-arm Doppler effect
In this section, we derive the single-arm Doppler effect in the GW space-based detectors induced by the axion-photon coupling, in the case where the OB modification shown in Fig. 3 is made. As a reminder, this coupling modifies the phase velocity of left and right circularly polarized light as derived in Eq. (8a), while the phase velocity of linearly polarized light is unchanged.
In Appendix. B, we derive the expected phase shift of light induced by vacuum birefringence on the detectors arms. It reads
$\Delta\phi(t)$ $=\frac{\sqrt{16\pi G\rho_{\mathrm{DM}}}E_{P}g_{a\gamma}}{\omega_{a}c}\sin\left(\frac{\omega_{a}L}{2c}\right)\sin\left(\omega_{a}t+\Phi^{\prime}\right)\,$ (9)
$\equiv\Delta\phi_{\mathrm{single\>link}}\,,$
for a right handed polarized EM wave where $\Phi^{\prime}=\Phi-\omega_{a}L/2c$. Therefore, following Fig. 4, we have
$\Delta\phi_{12}$ $=\Delta\phi_{23}=\Delta\phi_{31}=0\,$ (10a)
$\Delta\phi_{13}$ $=\Delta\phi_{32}=\Delta\phi_{21}=\Delta\phi_{\mathrm{single\>link}}\,,$ (10b)
where the $ij$ subscript denotes the arm pointing from S/C $j$ to S/C $i$. We define the relative frequency fluctuation signals (i.e Doppler shifts) as
$y_{ij}(t)$ $=\frac{1}{\omega_{j}}\frac{d\Delta\phi_{ij}}{dt}\,,$ (11)
where $\omega_{j}$ is the local laser frequency in S/C $j$. Using Eqs. (9), (10) and (11), we find the single-link Doppler shifts as
$y_{12}=y_{23}=y_{31}=0\,$ (12a)
$y_{13}(t)=y_{32}(t)=y_{21}(t)=\frac{\sqrt{16\pi G\rho_{\mathrm{DM}}}E_{P}g_{a\gamma}}{2\pi\nu_{0}c}\times\,$
$\sin\left(\frac{\omega_{a}L}{2c}\right)\cos\left(\omega_{a}t+\Phi^{\prime}\right)\,,$ (12b)
where $\omega_{j}/2\pi\approx\omega_{0}/2\pi=\nu_{0}$ is the laser frequency, supposed to be the same inside all OB.
IV.2 Sagnac $\alpha$ TDI combination
In space-based GW detectors, the laser noise on single link measurement is expected to dominate largely the signal by at least 8 orders of magnitude, which would make the GW detection impossible. Fortunately, there exists a method to effectively reduce drastically the laser noise by combining signals from the different S/C in such a way that the resulting observable is nearly independent of laser noise, but still contains the GW signal (or the DM signal in our case). This is called Time Delay Interferometry, or TDI and it was first introduced in Armstrong et al. (1999); Tinto and Armstrong (1999); Dhurandhar et al. (2002), for a detailed review see Tinto and Dhurandhar (2021). The principle relies on the fact that each laser noise is measured several times. Some specific time retarded signal combinations between the six different arms strongly reduces the laser noise to zero, thus the name of the method. In this section, we will only be interested in one specific TDI combination, the Sagnac $\alpha$, which is the most effective TDI combination to use for the axion-photon detection. Indeed, the Sagnac $\alpha$ measures the phase difference of two virtual laser beams starting at the same position and going around the constellation, one clockwise and the other counterclockwise. In our case, as Fig. 4 suggests, one of the two contributions (the one travelling clockwise) would get maximized, while the other one (the links travelling counterclockwise) would not contribute, and therefore the signal is maximum. In the following, we show how the one-link signal Eq. (12) can get amplified through the use of the Sagnac $\alpha$ combination.
In the time domain, the second generation Sagnac $\alpha$ combination is defined as Hartwig et al. (2023)
$\alpha(t)$ $=-\left(1-D_{1231}\right)\left[y_{13}+D_{13}y_{32}+D_{132}y_{21}-\right.\,$
$\left.\left(y_{12}+D_{12}y_{23}+D_{123}y_{31}\right)\right]\,,$ (13a)
where we use the compacted writing of TDI $D_{ij}y_{jk}(t)=y_{jk}\left(t-L_{ij}(t)/c\right)$ and $D_{ijk}y_{kl}(t)=y_{kl}\left(t-L_{ij}(t)/c-L_{jk}\left(t-L_{ij}(t)/c\right)\right)$. Using Eq. (12) and the constant and equal arm-length approximation (i.e. $L_{ij}(t)=L$), Eq. (13a) becomes
$\alpha(t)=\frac{\sqrt{16\pi G\rho_{\mathrm{DM}}}E_{P}g_{a\gamma}}{\pi\nu_{0}c}\sin^{2}\left(\frac{3\omega_{a}L}{2c}\right)\sin(\omega_{a}t+\Phi^{\prime\prime})\,,$ (13b)
with $\Phi^{\prime\prime}=\Phi^{\prime}-5\omega_{a}L/2c=\Phi-3\omega_{a}L/c$.
The one-sided power spectral density (PSD) of this signal reads
$S_{\alpha}(f)=$ $\left(\frac{\sqrt{16\pi G\rho_{\mathrm{DM}}}E_{P}g_{a\gamma}}{\pi\nu_{0}c}\right)^{2}\sin^{4}\left(\frac{3\omega_{a}L}{2c}\right)\,$
×$\quad\times\frac{1}{2\pi^{2}T_{\mathrm{obs}}}\left(\frac{\sin(\pi(f-f_{a})T_{\mathrm{obs}})}{f-f_{a}}\right)^{2}\,,$ (14)
where $f$ is the Fourier frequency, $f_{a}\equiv\omega_{a}/(2\pi)$ and $T_{\mathrm{obs}}$ is the total duration of the experiment. For simplicity we have used a square window of duration $T_{\mathrm{obs}}$ to calculate the PSD, which is sufficient for our rough sensitivity estimate.
The noise PSD of the $\alpha$ TDI channel depends mainly on two sources: (i) on the single-link optical metrology system (OMS) noise, which corresponds to the noise on the optical readout of the interference signals, and (ii) the single test mass acceleration noise, which corresponds to the noise coming from the inertial system comprising the test masses and all elements interacting directly with them. The shape and amplitude of those noises depend on the experiment under consideration. As a reminder, in this paper, we consider four space-based GW detectors: LISA, TianQin, Taiji and BBO. The first three of them consist in an individual triangular constellation either in a heliocentric orbit for LISA Colpi et al. (2024) and Taiji Luo et al. (2020) or geocentric one for TianQin Luo et al. (2016). BBO will operate four different triangular constellations, two of them located at almost the same position and the two other at different locations of the same heliocentric orbit Crowder and Cornish (2005).
In their current design, the PSD of the OMS and acceleration noises for LISA, TianQin and Taji are provided by
$S^{i}_{\mathrm{oms}}(f)$ $=\left(P^{i}_{\mathrm{oms}}\frac{2\pi f}{c}\right)^{2}\left(1+\left(\frac{2\>\mathrm{mHz}}{f}\right)^{4}\right)\>\mathrm{Hz}^{-1}\,$ (15a)
$S^{i}_{\mathrm{acc}}(f)$ $=\left(\frac{P^{i}_{\mathrm{acc}}}{2\pi fc}\right)^{2}\left(1+\left(\frac{0.4\>\mathrm{mHz}}{f}\right)^{2}\right)\,$
$\left(1+\left(\frac{f}{8\>\mathrm{mHz}}\right)^{4}\right)\>\mathrm{Hz}^{-1}\,,$ (15b)
[…]
$P^{\mathrm{LISA}}_{\mathrm{oms}}=15\times 10^{-12}\>\mathrm{m}\>,\>P^{\mathrm{LISA}}_{\mathrm{acc}}=3\times 10^{-15}\>\mathrm{m/s}^{2}\,$ (15c)
$P^{\mathrm{TianQin}}_{\mathrm{oms}}=10^{-12}\>\mathrm{m}\>,\>P^{\mathrm{TianQin}}_{\mathrm{acc}}=10^{-15}\>\mathrm{m/s}^{2}\,$ (15d)
$P^{\mathrm{Taiji}}_{\mathrm{oms}}=8\times 10^{-12}\>\mathrm{m}\>,\>P^{\mathrm{Taiji}}_{\mathrm{acc}}=3\times 10^{-15}\>\mathrm{m/s}^{2}\,,$ (15e)
while for BBO the single-link noises are given by Corbin and Cornish (2006)
$S^{\mathrm{BBO}}_{\mathrm{oms}}(f)$ $=\frac{1.0\times 10^{-34}}{(3L^{2})}\>\mathrm{m}^{2}/\mathrm{Hz}\,$ (15f)
$S^{\mathrm{BBO}}_{\mathrm{acc}}(f)$ $=\frac{4.5\times 10^{-34}}{(2\pi f)^{4}(3L)^{2}}\>\mathrm{m}^{2}.\mathrm{Hz}^{4}/\mathrm{Hz}\,,$ (15g)
where we added a factor $1/2$ to both contributions, to take into account the fact that the two triangular detectors located at the same position will probe the same amplitudes and phases of the field [^3] The two other detectors are much further away, and their distance is beyond one coherence length for frequencies $f\gtrapprox 1$ Hz, such that they will probe a different signal. Therefore we ignore them for our rough sensitivity estimate..
The modification of the OB suggested in Fig. 3 implies a decrease in the amplitude of light (or equivalently in the number of photons $N_{\gamma}$) that reaches each interferometer by a factor $2$ because light gets partially absorbed by the right polarizer. Since the OMS noise includes a contribution from shot noise whose PSD $\propto 1/N_{\gamma}$, we will conservatively add a factor $2$ to the whole OMS contribution [^4] The sensitivity to GW strain scaling as the square root of the TDI noise PSD, this implies a factor $\sqrt{2}$ loss of sensitivity on $h$., such that the Sagnac $\alpha$ noise PSD reads Hartwig et al. (2023) [^5] In Hartwig et al. (2023), only the first generation noise PSD are presented (see their Eq. (B2a)). In order to compute the second generation noise PSD for the Sagnac $\alpha$ combination, one needs to add an additional factor $(2\sin(3\omega L/2c))^{2}$.
$N_{\alpha}(f)$ $=4\sin^{2}\left(\frac{3\pi fL}{c}\right)\Bigg{[}12S_{\mathrm{oms}}(f)+$ (16)
$4\left(3-2\cos\left(\frac{2\pi fL}{c}\right)-\cos\left(\frac{6\pi fL}{c}\right)\right)S_{\mathrm{acc}}(f)\Bigg{]}\,.$
IV.3 Expected sensitivity
We now estimate the sensitivity to $g_{a\gamma}$ of each experiment under consideration in this paper. Since we assume a signal-to-noise ratio (SNR) of $1$, a simple (and conservative) estimate of the detectable coupling is obtained by equating the signal PSD $S_{\alpha}(f)$ (given by Eq. (14) in the limit $f\to f_{a}$) to the noise PSD $N_{\alpha}(f)$ (given by Eq. (16)), i.e.
$g_{a\gamma}(f)$ $=\frac{\pi\nu_{0}c}{\sqrt{8\pi G\rho_{\mathrm{DM}}}E_{P}\sin^{2}\left(\frac{3\pi fL}{c}\right)}\sqrt{\frac{N_{\alpha}(f)}{T_{\mathrm{obs}}}}\,.$ (17a)
Galactic DM is not purely monochromatic, it is a superposition of plane waves which oscillate at slightly different frequencies (see e.g. Derevianko (2018); Foster et al. (2018)). Therefore, the field acquires a characteristic coherence time $\tau(f)\sim 10^{6}/f$ Derevianko (2018). Eq. (17a) is valid when $T_{\mathrm{obs}}\ll\tau(f)$, i.e when the field can effectively be described as a monochromatic plane wave. Note that a correction factor to the sensitivity arises due to the stochastic nature of the amplitude of the field Centers et al. (2021) which induces a loss in signal of $\sim 1.5$ for a SNR 1. When the integration time is much longer than the coherence time of the field, i.e $T_{\mathrm{obs}}\gg\tau(f)$, the signal searched for is no longer coherent, i.e. it should be modeled as a sum of several stochastic harmonics, see Foster et al. (2018), and therefore Eq. (14) is not valid anymore. Another method to analyse the data is to cut the dataset in fragments with duration smaller than $\tau(f)$ and search for a coherent signal in each of these blocks of data. In such a case, the experimental sensitivity to the coupling is reduced and becomes Budker et al. (2014)
$g_{a\gamma}(f)$ $=\frac{\pi\nu_{0}c}{\sqrt{8\pi G\rho_{\mathrm{DM}}}E_{P}\sin^{2}\left(\frac{3\pi fL}{c}\right)}\sqrt{\frac{N_{\alpha}(f)}{\sqrt{T_{\mathrm{obs}}\tau(f)}}}\,.$ (17b)
{tblr}
vlines, colspec=ccccc LISA TianQin Taiji BBO
Laser frequency $\nu_{0}$ (Hz) $2.82\times 10^{14}$ $2.82\times 10^{14}$ $2.82\times 10^{14}$ $6\times 10^{14}$
Arm length $L$ (m) $2.5\times 10^{9}$ $\sqrt{3}\times 10^{8}$ $3\times 10^{9}$ $5\times 10^{7}$
Integration time $T_{\mathrm{obs}}$ (year) $4.5$ $5\times 1/2$ $5$ $4$
Frequency band (Hz) $[10^{-4}-1]$ $[10^{-4}-1]$ $[10^{-4}-1]$ $[10^{-1}-10]$
Table 1: Experimental parameters of interest for LISA, TianQin, Taiji and BBO, from Colpi et al. (2024); Quang Nam et al. (2023); Luo et al. (2016, 2020); Corbin and Cornish (2006); Cutler and Holz (2009). The time of integration of TianQin suffers from a $1/2$ factor because the detector plane is fixed throughout the orbit, implying it will be blinded by the Sun during half of the orbit. The time of integration of BBO is purely presumed.
All experimental parameters of interest for the sensitivity curves are extracted from Colpi et al. (2024); Quang Nam et al. (2023) for LISA, from Luo et al. (2016) for TianQin, from Luo et al. (2020) for Taiji and from Corbin and Cornish (2006); Cutler and Holz (2009) for BBO. They are summarized in Table  IV.3.
Figure 5: Sensitivity estimates of various GW space-based detectors (shown in brown, cyan, lime and magenta respectively for LISA, TianQin, Taiji and BBO) to the axion-photon coupling $g_{a\gamma}$. In light blue and light yellow, we show the respective constraint from CAST Anastassopoulos et al. (2017) and observations of SN1987A Payez et al. (2015) (from O’Hare (2020)). Finally, in dashed-dotted, dotted and dashed black lines, we show respectively the expected sensitivity of advanced LIGO, DECIGO and Cosmic Explorer detectors, from Nagano et al. (2019) (which have been rescaled for consistent numerical value of $\rho_{\mathrm{DM}}$).
In Fig. 5, we show the respective expected sensitivity of LISA, TianQin, Taiji and BBO, shown in brown, cyan, lime and magenta. One can notice that all these GW space-based detectors would improve the current laboratory constraints on $g_{a\gamma}$, for axion masses between $\sim 4\times 10^{-19}$ to $4\times 10^{-15}$ eV for LISA, TianQin and Taiji, and between $\sim 4\times 10^{-16}$ to $4\times 10^{-14}$ eV for BBO. In particular, between $10^{-17}$ and $10^{-15}$ eV, LISA, TianQin and Taiji would respectively reach $g_{a\gamma}\sim 10^{-12},1.5\times 10^{-12}$ and $4\times 10^{-13}$ GeV-1, improving by 1 order of magnitude the current best constraint at these masses, while BBO could reach $g_{a\gamma}\sim 10^{-16}$ GeV-1, 5 orders of magnitude better than SN1987A, the current best constraint at these masses.
V Discussion
In this paper, we have shown how operating circular polarized light beams in space-based gravitational wave detectors could transform them into axion detectors, whilst remaining sensitive to GW. Indeed, the coupling between axion and photon with strength $g_{a\gamma}$ induces an oscillation of the phase velocity of circularly polarized light through vacuum birefringence. In space-based GW detectors, it induces a phase shift of circularly polarized light compared to linearly polarized one, which would create an observable Doppler shift.
This could be made possible by slightly modifying the current OB design, changing one half waveplate by a circular polarizer. The number of photons that arrive at each optical interferometer from the distant S/C would be reduced by a factor 2, which would merely impact the minimum detectable GW amplitude by a small factor $\sqrt{2}$.
As we have shown it in Fig. 5, operating this new OB would allow the various space-based GW detectors under consideration in this paper to reach unconstrained regions of the $g_{a\gamma}-m_{a}$ parameter space, improving current best constraints on $g_{a\gamma}$ by as much as 5 orders of magnitude over several orders of magnitude of axion masses. In particular, such detection techniques would be complementary to futuristic versions of GW ground-based detectors, optimized for the axion-photon coupling searches Nagano et al. (2019), which are the most sensitive for axion masses between $\sim 10^{-16}$ and $10^{-11}$ eV (due to the smaller baseline).
Acknowledgements.
This work was supported by the Programme National GRAM of CNRS/INSU with INP and IN2P3 cofunded by CNES.
Appendix A OB modification
In this appendix, we show mathematically how the solution provided in Sec. III.4 is valid for both GW and axion detection.
A.1 Jones formalism
We can model the OB system using the Jones formalism for polarization states of light and optical elements which is very convenient when treating interference phenomena Collett (2005). This formalism discomposes the light polarization on a basis formed of the two linear polarizations. The polarization state is modelled using a $2\times 1$ matrix while the optical elements are represented by $2\times 2$ matrices.
As shown in Fig. 2, we assume the light beam to propagate along the $z$-axis with an initial polarization along the $x$-axis, that we call ”horizontal” polarization. Then, its Jones representation is Collett (2005)
$|H\rangle$ $=\begin{pmatrix}1\,\\ 0\end{pmatrix}\,.$ (18)
The other linear polarization along $y$-axis, that we will call ”vertical” is represented as
$|V\rangle$ $=\begin{pmatrix}0\,\\ 1\end{pmatrix}\,.$ (19)
A half-wave plate adds a phase $\phi=\pi$ to the light polarization parallel to its slow axis [^6] The principle of a wave plate is that it presents two perpendicular directions with respective refractive index $n_{f},n_{s}$ with $n_{f}<n_{s}$, that we call respectively fast and slow axis Hecht (2002). This leads to a phase shift between the light polarization along the fast axis and the one along the slow axis.. Its Jones matrix representation is Collett (2005)
$\hat{\lambda}_{2}(\theta)$ $=e^{-i\frac{\pi}{2}}\begin{pmatrix}\cos^{2}\theta-\sin^{2}\theta&2\sin\theta\cos\theta\,\\ 2\sin\theta\cos\theta&\sin^{2}\theta-\cos^{2}\theta\end{pmatrix}\,$
$\equiv\begin{pmatrix}\cos(2\theta)&\sin(2\theta)\,\\ \sin(2\theta)&-\cos(2\theta)\end{pmatrix}\,,$ (20)
up to an irrelevant total phase, where $\theta$ is the angle between the horizontal axis (the $x$-axis) and the fast axis of the plate. Similarly, a quarter-wave plate introduces a phase shift of $\pi/2$ between the two linear polarized states and can be represented as Collett (2005)
$\hat{\lambda}_{4}(\theta)$ $=\begin{pmatrix}\cos^{2}(\theta)-i\sin^{2}(\theta)&(1+i)\sin(\theta)\cos(\theta)\,\\ (1+i)\sin(\theta)\cos(\theta)&\sin^{2}(\theta)-i\cos^{2}(\theta)\end{pmatrix}\,,$ (21)
up to a total phase. Note that if the half or quarter wave plate is rotated such that its fast or slow axis is along the initial linear horizontal or vertical polarization of light (i.e $\theta=0$ or $\theta=\pi/2$), there is no effect, i.e the input and output polarization are the same (up to an irrelevant phase). Note also that the determinant of these matrices is $1$, i.e they preserve energy, or in other words, they have no absorption. In the Jones formalism, right and left circular polarizations are represented as Collett (2005)
$|R\rangle$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}1\,\\ i\end{pmatrix}\,$ (22a)
$|L\rangle$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}1\,\\ -i\end{pmatrix}\,.$ (22b)
We now introduce horizontal, vertical, left and right polarizers, which respectively transmit only horizontal, vertical, left and right polarizations of the input, defined as Fymat (1971)
$\hat{C}_{H}$ $=\begin{pmatrix}1&0\,\\ 0&0\end{pmatrix}\,$ (23a)
$\hat{C}_{V}$ $=\begin{pmatrix}0&0\,\\ 0&1\end{pmatrix}\,$ (23b)
$\hat{C}_{L}$ $=\frac{1}{2}\begin{pmatrix}1&i\,\\ -i&1\end{pmatrix}\,$ (23c)
$\hat{C}_{R}$ $=\frac{1}{2}\begin{pmatrix}1&-i\,\\ i&1\end{pmatrix}\,.$ (23d)
In practice, the right polarizer is made of the combination $\hat{\lambda}_{4}(\pi/4)\hat{C}_{H}\hat{\lambda}_{4}(-\pi/4)$ (and for the left polarizer, one replaces $\hat{C}_{H}$ by $\hat{C}_{V}$ or exchange the two quarter-wave plates). Finally, a mirror reverses the direction of one of the two linear polarizations while letting the other one unchanged. Its representation is Fymat (1971)
$\hat{M}=\begin{pmatrix}-1&0\,\\ 0&1\end{pmatrix}\,,$ (24)
therefore it is the ”horizontal” polarization that gets flipped.
A.2 Requirements for GW and axion detection
Between the polarized beam splitter and the end of the OB, we will model the set of optical elements of $n$ half-wave plates, $m$ quarter-wave plates and $\ell$ polarizers (of any kind) that we add by $\hat{S}(\theta,\theta^{\prime})$, where $\theta(\theta^{\prime})$ is the angle between the fast axis of the half-wave (quarter-wave) plates and the horizontal axis. We will denote this set of elements with superscript $A$ or $B$ for the OB $A$ or $B$, which are respectively associated with angles $\theta_{A}(\theta_{B}),\theta^{\prime}_{A}(\theta^{\prime}_{B})$, i.e we note $\hat{S}^{(A)}(\theta_{A},\theta^{\prime}_{A})$ and $\hat{S}^{(B)}(\theta_{B},\theta^{\prime}_{B})$.
The first requirement (noted $R_{1}$) is that the incoming polarization from space can interfere with the local polarization, and is not fully rejected by the polarized beam splitter. For an initial $|H\rangle$ polarization state for the beam on both S/C, the polarization of the beam after leaving the OB $A$ and $B$ is respectively $\hat{S}^{(A)}(\theta_{A},\theta^{\prime}_{A})|H\rangle$ and $\hat{S}^{(B)}(\theta_{B},\theta^{\prime}_{B})|H\rangle$. Then, after travelling in free space, each light beam enters the other S/C and goes through the set of optical elements but with opposite propagation direction. In such a case, light sees the fast axis with opposite angle compared to horizontal plane such that the total polarization of the beam becomes respectively
$\hat{S}^{(B)}(-\theta_{B},-\theta^{\prime}_{B})\hat{S}^{(A)}(\theta_{A},\theta^{\prime}_{A})$ $|H\rangle\,$ (25a)
$\hat{S}^{(A)}(-\theta_{A},-\theta^{\prime}_{A})\hat{S}^{(B)}(\theta_{B},\theta^{\prime}_{B})$ $|H\rangle\,.$ (25b)
$R_{1}$ implies that both polarizations Eq. (25a) are not a pure ”horizontal” linear polarization, i.e
$R_{1}\Rightarrow\left\{\begin{array}[]{rcl}\hat{S}^{(B)}(-\theta_{B},-\theta^{\prime}_{B})\hat{S}^{(A)}(\theta_{A},\theta^{\prime}_{A})|H\rangle&\neq|H\rangle\\ \hat{S}^{(A)}(-\theta_{A},-\theta^{\prime}_{A})\hat{S}^{(B)}(\theta_{B},\theta^{\prime}_{B})|H\rangle&\neq|H\rangle\,,\end{array}\right.$ (28)
up to a constant phase.
The second requirement ($R_{2}$) for the optical interferometer to work is that the reflected component of light is the same as the initial one, such that it is rejected by the beam splitter, i.e we require that after travelling inside the set $\hat{S}(\theta,\theta^{\prime})$, getting reflected on the mirror and travelling back inside $\hat{S}(\theta,\theta^{\prime})$ (but with opposite propagation direction, as before), the polarization is ”horizontal”. This means
$R_{2}\Rightarrow\left\{\begin{array}[]{rcl}\hat{S}^{(A)}(-\theta_{A},-\theta^{\prime}_{A})\hat{M}S^{(A)}(\theta_{A},\theta^{\prime}_{A})|H\rangle&=|H\rangle\\ \hat{S}^{(B)}(-\theta_{B},-\theta^{\prime}_{B})\hat{M}S^{(B)}(\theta_{B},\theta^{\prime}_{B})|H\rangle&=|H\rangle\,,\end{array}\right.$ (31)
still up to a constant phase.
Finally, the additional requirement ($R_{3}$) to be sensitive to the axion field is that at least one of the two light polarizations in outer space has to be circularly polarized. If we assume the light propagating clockwise is right circularly polarized, then, the light propagating counterclockwise can be anything else, i.e linearly polarized or left polarized state [^7] If it is linearly polarized or left polarized, the beatnote between the two beams will oscillate at the axion frequency, but the amplitude will be twice as big in the second case.. Mathematically,
$R_{3}\Rightarrow\left\{\begin{array}[]{rcl}\hat{S}^{(A)}(\theta_{A},\theta^{\prime}_{A})|H\rangle&=|R\rangle\\ \hat{S}^{(B)}(\theta_{B},\theta^{\prime}_{B})|H\rangle&\neq|R\rangle\,,\end{array}\right.$ (34)
up to a constant phase.
A.3 Solution
As mentioned in the main text, a solution to all the requirements of the last section is to replace the last $\hat{\lambda}_{2}$ in the OB $A$ by a right polarizer i.e
$\hat{S}^{(A)}$ $=\hat{C}_{R}\,$ (35a)
$\hat{S}^{(B)}$ $=\mathds{1}_{2}\,,$ (35b)
where the latter is the identity matrix We show explicitly how it resolves the various requirements. Starting with $R_{3}$, we have
$\hat{S}^{(A)}|H\rangle$ $=|R\rangle\,$ (36a)
$\hat{S}^{(B)}|H\rangle$ $=|H\rangle\neq|R\rangle\,,$ (36b)
and therefore, in free space, one polarization is right circularly polarized and the other linearly polarized such that the axion field impacts only one of the two beams (see Fig. 3).
Then, looking at $R_{1}$, we have
$\hat{S}^{(B)}\hat{S}^{(A)}|H\rangle$ $=|R\rangle\neq|H\rangle\,$ (37a)
$\hat{S}^{(A)}\hat{S}^{(B)}|H\rangle$ $=|R\rangle\neq|H\rangle\,,$ (37b)
i.e the light coming from the OB $A$ can interfere in the interferometer in OB $B$ and vice-versa. Indeed, the light leaving OB $A$ is right circularly polarized, arrives in OB $B$ and since a circular polarization is a linear superposition of the two linear polarizations, the vertical component is reflected towards the interferometer, as shown in Fig. 3. The other beam coming from OB $B$ is in the ”horizontal” polarization state in vacuum, arrives on the right polarizer of OB $A$, gets right circularly polarized and similarly as above, its vertical component gets reflected by the polarized beam splitter inside the interferometer (see Fig. 3).
Finally, let us make sure that no parasitic component of light gets inside the interferometers after getting reflected at the edge of the S/C (requirement $R_{2}$). First, as we have not modified anything in the OB $B$, we have
$\hat{S}^{(B)}\hat{M}\hat{S}^{(B)}|H\rangle$ $=|H\rangle\,,$ (38)
i.e the reflected component is in the ”horizontal” polarization state, and therefore transmitted by the polarized beam splitter. Then, it is easy to show that
$\hat{C}_{R}\hat{M}\hat{C}_{R}$ $=0\,,$ (39a)
such that
$\hat{S}^{(A)}\hat{M}\hat{S}^{(A)}|H\rangle$ $=0\,,$ (39b)
i.e the entire power reflected at the edge of OB $A$ is absorbed by the right polarizer and therefore does not jeopardize the functioning of the interferometer. This can be intuitively understood by the fact that after reflection, the right circularly polarization becomes left circularly polarized and goes back inside the right polarizer, whose role is to extract and transmit only the right polarization of the input. Therefore, nothing from the left polarized parasitic light goes through.
Appendix B Phase shift induced by vacuum birefringence
In this appendix, we derive the phase shift induced by vacuum birefringence on the detector arms. The axion-photon coupling affects the optical length of the arm via the light travel time $\tau_{\rightarrow}(t)$ (expressed as a function of the receiving time $t$) from one S/C to the other as
[…]
where $c(t)$ is the time dependent phase velocity of light whose exact form depends on the light polarization Eq. (8a), the sign of $\pm$ depends on the right ($+$) or left (--) polarization and $\delta c$ is given by Eq. (8c). To first order in $\delta c$, the previous equation becomes
[…]
depending on the right or left polarization.
Assuming a right handed photon (i.e. taking the $+$ sign in the previous equation), the photon travel time (as a function of the receiving time) is given by
[…]
We can then interpret this oscillation of phase velocity of light as an oscillation of the distance between the S/C itself, i.e we can write
$\tau_{\rightarrow}(t)$ $=\frac{L-\delta L(t)}{c}\,,$ (43a)
where, using Eq. (8c)
[…]
$=\frac{\sqrt{16\pi G\rho_{\mathrm{DM}}}E_{P}g_{a\gamma}}{\omega_{a}kc}\sin\left(\frac{\omega_{a}L}{2c}\right)\sin\left(\omega_{a}\left(t-\frac{L}{2c}\right)+\Phi\right)$ (43c)
Compared to a situation where $\delta L=0$ (typically for linearly polarized light), the phase accumulated by light due to this armlength variation is simply
$\Delta\phi(t)=k\>\delta L(t)\,$
$=\frac{\sqrt{16\pi G\rho_{\mathrm{DM}}}E_{P}g_{a\gamma}}{\omega_{a}c}\sin\left(\frac{\omega_{a}L}{2c}\right)\sin\left(\omega_{a}\left(t-\frac{L}{2c}\right)+\Phi\right)\,.$ (44)